# Machine-directed gravitational-wave counterpart discovery

### v2

@@ 1 Introduction
The inspiral and subsequent merger of compact objects emits gravitational waves (GWs) that can be detected by ground-based interferometers like LIGO (LSC et al., 2015), Virgo (Acernese et al., 2014), and KAGRA (Akutsu et al., 2021) (also the International Gravitational-wave Network; IGWN). If at least one of the binary components is a neutron star (NS), these phenomena may have analogs in electromagnetic (EM) waves with signatures spanning the EM spectrum, from gamma rays to radio waves. Kilonovae (KNe) are ultraviolet, optical, near-infrared transients that arise due to rapid neutron capture (r-process) nucleosynthesis in the merger ejecta (Lattimer & Schramm, 1974; Symbalisty & Schramm, 1982; Li & Paczyński, 1998). They are robust counterparts to most binary neutron star (BNS) and many neutron-star black hole (NSBH) mergers (Metzger et al., 2010). Their light curves and spectra provide the opportunity to probe the ejecta and processes driving nucleosynthesis in it (Kasen et al., 2017; Drout et al., 2017; Kasliwal et al., 2017), and hypermassive/supramassive NS formation scenarios to place independent constraints on the NS equation of state (EoS) (Bauswein et al., 2017; Radice et al., 2018). However, they are quite elusive because they are faint (Mless-$\lesssim$ -16 mag in optical) and short-lived ($\lesssim$1 week) (Bulla, 2019; Metzger, 2019).
@@
AT2017gfo was the first, and so far only, kilonova associated with a GW signal (GW170817, Abbott et al., 2017) discovered (Coulter et al., 2017; Smartt et al., 2017). This milestone was achieved due to data in all messengers being broadcast worldwide in real-time and cyberinfrastructure/algorithms for unifying and characterizing them. On the EM side, this includes General Coordinates Network (GCN) notices of gamma ray bursts (GRBs) associated with jets launched by accretion onto the merger remnant (Narayan et al., 1992), new transients discovered by ground-based surveys (Kochanek et al., 2017; Tonry et al., 2018; Bellm, 2014; Bellm & Kulkarni, 2017) broadcast via Astronote/Transient Name Server and data brokers (Förster et al., 2021; Nordin et al., 2019; Saha et al., 2014; Möller et al., 2021), and follow-up observations co-ordinated by marshals/Target and Observation Managers (Street et al., 2018; Coughlin et al., 2023) and performed by networks of telescopes (Antier et al., 2020; Kasliwal et al., 2020). While each of these solutions represent important components of the required cyberinfrastructure for identifying KNe, they are not sufficient.
@@
When the IGWN detects a signal from a compact binary coalescence, they issue a GCN notice to the astronomy community, along with some GW derived inference, like source sky localization and distance [^1] https://emfollow.docs.ligo.org/userguide/. Depending on the orientation of the system, satellites could detect a gamma-ray burst (GRB) powered by the merger and help reduce the search area by crossmatching credible regions. In fact, several GRB afterglow target-of-opportunity (ToO) searches (Singer et al., 2015; Ahumada et al., 2022) have paved the way for GW ToO campaigns (Kasliwal et al., 2020). Transients that are temporally and spatially coincident with the GW event are assembled. The vast majority of contaminants in this list are supernovae, cataclysmic variables (CVs), and unassociated GRB afterglows. Especially at early phases, these can be difficult to distinguish from KNe. Human experts manually examine the candidate light curves together with additional information, like host galaxies/environment, preliminary fits to light curve models (e.g. kilonovae and common imposters, such as shock cooling emission), and perform additional follow-up to obtain better characterization, and hopefully identify the real event. This step can occur at late hours via oral/written electronic communication and is the one of the main sources of inefficiency in GW counterpart searches.
@@
Optimal resource allocation has received a lot of interest given the deluge of survey data and limited follow-up resources. To first order this can be performed using population analyses (e.g., Carbone & Corsi, 2018; Williamson et al., 2019; Carbone & Corsi, 2020) to identify strategies that optimize a given metric for the entire population instead of being customized to each event. More targeted approaches, e.g. event selection for spectroscopic follow-up, have recently been developed using sophisticated machine learning approaches (Ishida et al., 2019; Astudillo et al., 2020; Kennamer et al., 2020; Wang & Melchior, 2022). RL was recently used to optimize spectroscopic follow-up of galaxies to maximize Shannon Information on cosmological parameters (Cranmer et al., 2021). In contrast, our work focuses on real-time resource allocation. The use of RL to plan the follow-up of transients is necessary because information gained from follow-up may only be apparent after all the data arrives (discussed in Section 2). RL agents are also robust to stochasticity, which is intrinsic to survey data and the execution of ToOs. These inherent complexities render approaches such as linear programming or rule-based systems inadequate.
@@ 2 Autonomous Control
To solve sequential decision making under uncertainty problems one needs to consider caveats that make it very different from vanilla machine learning. First, there is no longer a static distribution of data (assumed identically and independently distribute or i.i.d) represented by features $X$, and in supervised learning, target $Y$, where the goal is the learn a generalized function $f(X)$, that in supervised learning maps to $Y$. The key issue in adaptive control problems is that the distribution of observed data evolves as a result of the AI agent interacting with it and is no longer i.i.d. This is because the actions chosen by the agent determine the next set of observations, making them highly correlated. Using supervised learning in these situations can lead to compounding errors as the data distribution starts deviating from an initial dataset (Ross et al., 2011). Closely related is the idea of exploration. In vanilla machine learning, one assumes that the training data is all there is and the prediction task will be on a dataset that follows a similar distribution. There is no notion of acquiring more data to improve on the optimization problem. An exception is active learning, where labels from an oracle (e.g. a human expert) are acquired to improve the algorithm’s performance. However, even in this case, the training and target data distribution are assumed to remain i.i.d. Another key difference is the notion of delayed consequences. Using the KN follow-up example, an agent may take a decision to obtain follow-up today but it will not be clear until the entire data comes as to how important the resulting data (and therefore how good the decision) was for constraining light curve physics. This makes it hard to solve the so-called “credit assignment problem”, which is to identify which action among a series of actions produced the desirable outcome.
@@
The state is a mathematical representation of the world that the agent observes. For the KN follow-up problem this is data contained in GCNs, survey data (light curves and image stamps), and any miscellaneous information relevant to the problem. The action is a mathematical representation of how an agent responds. For the KN photometric follow-up problem, this might be choosing among a set of follow-up observations, e.g. {no action, observe $g$-band, observe $r$-band, observe $g$-band & $r$-band}. Actions may be associated with costs and subject to a budget. The policy is a function $\pi(a\mid s)=P(a_{t}=a\mid s_{t}=s)$, mapping the agent’s observed state to a distribution of actions it takes. The reward is the utility of taking a given action in a given state. This encodes the science objective needing maximization. For the KN follow-up case, this could be the improvement in constraints on model parameters. Such rewards will be delayed, making the use of RL indispensable for such applications.
In an MDP, at timestep $t$, an agent observes a state of the world $s_{t}$ returned by the environment and takes an action $a_{t}$ according to its policy. The environment returns to the agent a new state $s_{t+1}$ along with a reward $r_{t}$ and the process repeats, either forever or until a horizon, $H$, is reached.
In many RL problems, one requires the Markov property be satisfied, in that the state representation is a sufficient statistic of the full history. Specifically,
@@
For finite states, the idea is to estimate $Q$ as an iterative update using the Bellman identity. Due to the contraction property of the Bellman operator, such policy iteration algorithms converge to the optimal action-value function, $Q^{*}$, as $i\to\infty$, where $i$ is the iteration count. In most RL problems, it is computationally infeasible to compute the expectation over next states (denoted by the sum in Equation 2). Therefore, it is common to sample the expectation using Monte Carlo or Temporal Difference (TD) methods (Tesauro, 1995). Notably, TD methods both sample and bootstrap by approximating the expectation as $Q(s_{t+1},a_{t+1})$ in SARSA (Rummery & Niranjan, 1994) or $\max_{a}Q(s_{t+1},a)$ in Q-learning (Watkins, 1989), where Q is the bootstrapped using the running estimate. TD estimates are biased (as it bootstraps) and high variance. However, TD methods are more versatile than Monte Carlo as they can handle infinite horizon MDPs. In this work we use TD methods.
In most realistic situations, where the state space is continuous, such an iterative update is not computationally tractable since the cardinality of $Q(s,a)$ is infinite. In practice, it is therefore common to use a function approximator to estimate $Q(s,a;\theta)$, where $\theta$ are parameters describing $Q$. In RL, traditionally this is a linear value function approximator (VFA, Tesauro, 1995), where $Q$ is parameterized by linear weights ($w$) as a function of state-action features, i.e. $Q(s,a;w)=x(s,a)^{T}w$. More recently, non-linear function approximators (like neural networks) have been used (Mnih et al., 2013) with much success. Updates to $Q$ are then achieved using stochastic gradient descent along the mean squared error between the Q approximation (e.g. TD estimate) and the running estimate. In this work, we adopt linear VFA in SARSA despite its simplicity because the algorithm is guaranteed to converge to within a constant factor of the optimal behavior policy under minor assumptions (Tsitsiklis & Van Roy, 1997, discussed in detail in Section 3.3). This is not the case for Q-learning or non-linear approximators. This attribute makes it suitable for our pilot investigation where we are interesting exploring the suitability of RL approaches to solve the KN follow-up problem and providing useful benchmarks on the same.
@@ 3.1 Problem Statement
Our agent is presented $N$ transient light curves from ZTF, one of which is the KN and the rest are contaminants, chosen randomly from a list of supernovae and unassociated GRB afterglows. The agent observes the candidates on day 1. On days 2 through 7 it assigns one additional photometry point with ZTF in $g$, $r$, or $i$ using a deep 300s exposure to one of the events. The agent gets a reward 1 if the follow-up is assigned to the KN, and 0 otherwise. The objective is learn a policy to maximize the number of follow-up assigned to the true KN. The maximum achievable score is 6. A random agent will achieve an expected score of $6/N$.
We have chosen a non-model specific reward that seeks to maximize sampling, with that expectation that more data leads to better constraints. KNe are quite novel and there are still a lot of open questions surrounding ejecta composition and the treatment of opacities and its effect on the KN light curve (Metzger, 2019). Targeting constraining models could then result in poor generalizability or transfer to real-world domains of learned behavior policies. However, given reliable and representative models, constraining physics parameters could be a good choice.
@@ 3.2 Training Data
We generate training light curves as follows. We place transients randomly in the 3-D volume of the Universe observable by ZTF. ZTF is a public-private partnership survey of the Northern sky conducted in $g$, $r$, and $i$ filters using a wide field (47 deg2) imager on the Palomar 48-inch telescope (Masci et al., 2019; Dekany et al., 2020; Graham et al., 2019; Bellm et al., 2019). For Type Ia supernovae we use SALT2 (Guy et al., 2007) models to estimate photometry in ZTF filters. We apply Milky way extinction (Cardelli et al., 1989; Schlegel et al., 1998) given the supernova sky location, redshift and K-corrections. For core-collapse supernovae, we use 2-D Gaussian Process fits to ZTF light curves of spectroscopically confirmed supernovae of types II, IIn, IIb, Ib, Ic, Ic-BL, and superluminous supernovae types I and II. While we apply redshift corrections (scaling and stretch), due the empirical nature of the fit, we are unable to apply extinction corrections. We estimate K-correction by estimating the Gaussian Process fit in the wavelength dimension at the redshifted mean wavelength of the ZTF filter at the sampled redshift. For GRB afterglows, we use the semi-analytic models of Ryan et al. (2020). Finally, we use the models from Bulla (2019) to simulate KNe from BNS and NSBH mergers. For KNe we use the interpolation scheme in Pang et al. (2023). For these, once again, while redshift scaling is accounted for, we do not account for extinction or K-correction.
@@
Figure 1: Learning curve of our KN follow-up agent Pythia (blue). Our agent achieves >3.5×>3.5$\times$> 3.5 × higher score than a random agent during training. The score variance (shaded blue region) is computed using a moving average with a length of 50 episodes. The mean random score is 6/9. Human references are show as orange shaded regions.
@@ 3.3 RL Algorithm
We learn the optimal behavior policy using SARSA. SARSA is a simple on-policy learning algorithm, i.e. samples from the policy are used to update $Q$ (see Algorithm 2). We parametrize $Q=x(s,a)^{T}w$, where $x(s,a)$ is the state-action feature representation and $w$ is a set of linear weights to be learned. We choose a TD(0) target, i.e. $r_{t}+\gamma\hat{Q}(s_{t+1},a_{t+1};\hat{w})$, where $\hat{Q}$ is the running estimate. The policy is $\epsilon$-greedy: with probability $\epsilon$ the agent chooses a random action and otherwise chooses $\max_{a}\hat{Q}(s_{t},a_{t};\hat{w})$, the action with the maximum $\hat{Q}$-value. If the learning rate $\alpha_{t}$ satisfies the Robbins-Munro sequence and $\epsilon\to 0$ as $i\to\infty$ the agent is Greedy in the Limit of Infinite Exploration (GLIE) and is a sufficient condition for convergence in SARSA [^3] Two conditions to be GLIE: (1) we visit every state-action pair infinite times and (2) behavior policy converges to greedy policy. We parameterize $\epsilon=1/k^{n}$, where $n$ is a hyperparameter that controls how fast the agent becomes greedy. This parametrization allows the agent to get in a lot of exploration early during training and gradually forces it become more greedy. To ensure sample efficiency and propagate TD updates faster, we reuse episodes $m$ times, where we set $m=5$. Our agent learns online (i.e. it collects new experiences as opposed to episodes collected beforehand) in a simulated environment [^4] It is possible to train the agent online with real GW trigger scenarios but RL algorithms are notoriously sample inefficient and can take a lot of exploration to arrive at the optimal $Q$/policy..
@@
Our choice of linear VFA was motivated by the fact that the triad of non-linear function approximators, off-policy learning (like Q-learning), and bootstrapping (as in Temporal-difference learning) can cause the $Q$ function to diverge (Tsitsiklis & Van Roy, 1997). Though recent works have made significant progress to address this issue (e.g. prioritized reply (Schaul et al., 2015), asynchronous methods (Mnih et al., 2016), dueling DQN targets (Wang et al., 2016)) we adopt linear VFA and SARSA for our toy implementation to assess the feasibility of RL approaches to solve the KN follow-up problem and disambiguate limitations from convergence issues. An issue may be that the choice of linear VFA as hypothesis function class may not be a sufficiently rich representation of the true $Q$ function. This does indeed appear to be the case as we discuss later.

### v3

@@ 1 Introduction
The inspiral and subsequent merger of compact objects emits gravitational waves (GWs) that can be detected by ground-based interferometers like LIGO (LSC et al., 2015), Virgo (Acernese et al., 2014), and KAGRA (Akutsu et al., 2021) (also the International Gravitational-wave Network; IGWN). If at least one of the binary components is a neutron star (NS), these phenomena may have analogs in electromagnetic (EM) waves with signatures spanning the EM spectrum, from gamma rays to radio waves. Kilonovae (KNe) are ultraviolet, optical, near-infrared transients that arise due to rapid neutron capture (r-process) nucleosynthesis in the merger ejecta (Lattimer & Schramm, 1974; Symbalisty & Schramm, 1982; Li & Paczyński, 1998). They are robust counterparts to most binary neutron star (BNS) and many neutron-star black hole (NSBH) mergers (Metzger et al., 2010). Their light curves and spectra provide the opportunity to probe the ejecta and processes driving nucleosynthesis in it (Kasen et al., 2017; Drout et al., 2017; Kasliwal et al., 2017), and hypermassive/supramassive NS formation scenarios to place independent constraints on the NS equation of state (EoS) (Bauswein et al., 2017; Radice et al., 2018). However, they are quite elusive because they are faint (Mless-$\lesssim$ -16 mag in optical) and short lived ($\lesssim$1 week) (Bulla, 2019; Metzger, 2019).
@@
AT2017gfo was the first, and so far only, kilonova associated with a GW signal (GW170817, Abbott et al., 2017) discovered (Coulter et al., 2017; Smartt et al., 2017). This milestone was achieved due to data in all messengers being broadcast worldwide in real-time and cyberinfrastructure/algorithms for unifying and characterizing them. On the EM side, this includes General Coordinates Network (GCN) notices of gamma ray bursts (GRBs) associated with jets launched by accretion onto the merger remnant (Narayan et al., 1992), new transients discovered by ground-based surveys (Kochanek et al., 2017; Tonry et al., 2018; Bellm, 2014; Bellm & Kulkarni, 2017) broadcast via Astronote/Transient Name Server and data brokers (Förster et al., 2021; Nordin et al., 2019; Saha et al., 2014; Möller et al., 2021), and follow-up observations co-ordinated by marshals/Target and Observation Managers (Street et al., 2018; Coughlin et al., 2023) and performed by networks of telescopes (Antier et al., 2020; Kasliwal et al., 2020). While each of these solutions represents important components of the required cyberinfrastructure for identifying KNe, they are not sufficient.
@@
When the IGWN detects a signal from a compact binary coalescence, they issue a GCN notice to the astronomy community, along with some GW-derived inference, like source sky localization and distance [^1] https://emfollow.docs.ligo.org/userguide/. Depending on the orientation of the system, satellites could detect a gamma-ray burst (GRB) powered by the merger and help reduce the search area by crossmatching credible regions. In fact, several GRB afterglow target-of-opportunity (ToO) searches (Singer et al., 2015; Ahumada et al., 2022) have paved the way for GW ToO campaigns (Kasliwal et al., 2020). Transients that are temporally and spatially coincident with the GW event are assembled. The vast majority of contaminants in this list are supernovae, cataclysmic variables (CVs), and unassociated GRB afterglows. Especially at early phases, these can be difficult to distinguish from KNe. Human experts manually examine the candidate light curves together with additional information, like host galaxies/environment, preliminary fits to light curve models (e.g. kilonovae and common imposters, such as shock cooling emission), and perform additional follow-up to obtain better characterization, and hopefully identify the real event. This step can occur at late hours via oral/written electronic communication and is the one of the main sources of inefficiency in GW counterpart searches.
@@
Optimal resource allocation has received a lot of interest given the deluge of survey data and limited follow-up resources. To first order, this can be performed using population analyses (e.g., Carbone & Corsi, 2018; Williamson et al., 2019; Carbone & Corsi, 2020) to identify strategies that optimize a given metric for the entire population instead of being customized to each event. More targeted approaches, e.g. event selection for spectroscopic follow-up, have recently been developed using sophisticated machine learning approaches (Ishida et al., 2019; Astudillo et al., 2020; Kennamer et al., 2020; Wang & Melchior, 2022). Reinforcement Learning (RL) was recently used to optimize spectroscopic follow-up of galaxies to maximize Shannon Information on cosmological parameters (Cranmer et al., 2021). In contrast, our work focuses on real-time resource allocation. The use of RL to plan the follow-up of transients is necessary because information gained from follow-up may only be apparent after all the data arrives (discussed in Section 2). RL agents are also robust to stochasticity, which is intrinsic to survey data and the execution of ToOs. These inherent complexities render approaches such as linear programming or rule-based systems inadequate.
@@ 2 Autonomous Control
To solve sequential decision making under uncertainty problems one needs to consider caveats that make it very different from vanilla machine learning. First, there is no longer a static distribution of data (assumed identically and independently distribute or i.i.d) represented by features $X$, and in supervised learning, target $Y$, where the goal is to the learn a generalized function $f(X)$, that in supervised learning maps to $Y$. The key issue in adaptive control problems is that the distribution of observed data evolves as a result of the AI agent interacting with it and is no longer i.i.d. This is because the actions chosen by the agent determine the next set of observations, making them highly correlated. Using supervised learning in these situations can lead to compounding errors as the data distribution starts deviating from an initial dataset (Ross et al., 2011). Closely related is the idea of exploration. In vanilla machine learning, one assumes that the training data is all there is and the prediction task will be on a dataset that follows a similar distribution. There is no notion of acquiring more data to improve on the optimization problem. An exception is active learning, where labels from an oracle (e.g. a human expert) are acquired to improve the algorithm’s performance. However, even in this case, the training and target data distribution are assumed to remain i.i.d. Another key difference is the notion of delayed consequences. Using the KN follow-up example, an agent may take a decision to obtain follow-up today but it will not be clear until the entire data comes as to how important the resulting data (and therefore how good the decision) was for constraining light curve physics. This makes it hard to solve the so-called “credit assignment problem”, which is to identify which action among a series of actions produced the desirable outcome.
@@
The state is a mathematical representation of the world that the agent observes. For the KN follow-up problem, this is data contained in GCNs, survey data (light curves and image stamps), and any miscellaneous information relevant to the problem. The action is a mathematical representation of how an agent responds. For the KN photometric follow-up problem, this might be choosing among a set of follow-up observations, e.g. {no action, observe $g$-band, observe $r$-band, observe $g$-band & $r$-band}. Actions may be associated with costs and subject to a budget. The policy is a function $\pi(a\mid s)=P(a_{t}=a\mid s_{t}=s)$, mapping the agent’s observed state to a distribution of actions it takes. The reward is the utility of taking a given action in a given state. This encodes the science objective needing maximization. For the KN follow-up case, this could be the improvement in constraints on model parameters. Such rewards will be delayed, making the use of RL indispensable for such applications.
In an MDP, at timestep $t$, an agent observes the state of the world $s_{t}$ returned by the environment and takes an action $a_{t}$ according to its policy. The environment returns to the agent a new state $s_{t+1}$ along with a reward $r_{t}$ and the process repeats, either forever or until a horizon, $H$, is reached.
In many RL problems, one requires the Markov property to be satisfied, in that the state representation is a sufficient statistic of the full history. Specifically,
@@
For finite states, the idea is to estimate $Q$ as an iterative update using the Bellman identity. Due to the contraction property of the Bellman operator, such policy iteration algorithms converge to the optimal action-value function, $Q^{*}$, as $i\to\infty$, where $i$ is the iteration count. In most RL problems, it is computationally infeasible to compute the expectation over next states (denoted by the sum in Equation 2). Therefore, it is common to sample the expectation using Monte Carlo or Temporal Difference (TD) methods (Tesauro, 1995). Notably, TD methods both sample and bootstrap by approximating the expectation as $Q(s_{t+1},a_{t+1})$ in SARSA (Rummery & Niranjan, 1994) or $\max_{a}Q(s_{t+1},a)$ in Q-learning (Watkins, 1989), where Q is the bootstrapped using the running estimate. TD estimates are biased (as it bootstraps) and have high variance. However, TD methods are more versatile than Monte Carlo as they can handle infinite horizon MDPs. In this work, we use TD methods.
In most realistic situations, where the state space is continuous, such an iterative update is not computationally tractable since the cardinality of $Q(s,a)$ is infinite. In practice, it is therefore common to use a function approximator to estimate $Q(s,a;\theta)$, where $\theta$ are parameters describing $Q$. In RL, traditionally this is a linear value function approximator (VFA, Tesauro, 1995), where $Q$ is parameterized by linear weights ($w$) as a function of state-action features, i.e. $Q(s,a;w)=x(s,a)^{T}w$. More recently, non-linear function approximators (like neural networks) have been used (Mnih et al., 2013) with much success. Updates to $Q$ are then achieved using stochastic gradient descent along the mean squared error between the Q approximation (e.g. TD estimate) and the running estimate. In this work, we adopt linear VFA in SARSA despite its simplicity because the algorithm is guaranteed to converge to within a constant factor of the optimal behavior policy under minor assumptions (Tsitsiklis & Van Roy, 1997, discussed in detail in Section 3.3). This is not the case for Q-learning or non-linear approximators. This attribute makes it suitable for our pilot investigation where we are interested in exploring the suitability of RL approaches to solve the KN follow-up problem and providing useful benchmarks on the same.
@@ 3.1 Problem Statement
Our agent is presented $N$ transient light curves from ZTF, one of which is the KN and the rest are contaminants, chosen randomly from a list of supernovae and unassociated GRB afterglows. The agent observes the candidates on day 1. On days 2 through 7 it assigns one additional photometry point with ZTF in $g$, $r$, or $i$ using a deep 300s exposure to one of the events. The agent gets a reward 1 if the follow-up is assigned to the KN, and 0 otherwise. The objective is to learn a policy to maximize the number of follow-up assigned to the true KN. The maximum achievable score is 6. A random agent will achieve an expected score of $6/N$.
We have chosen a non-model specific reward that seeks to maximize sampling, with the expectation that more data leads to better constraints. KNe are quite novel and there are still a lot of open questions surrounding ejecta composition and the treatment of opacities and its effect on the KN light curve (Metzger, 2019). Targeting constraining models could then result in poor generalizability or transfer to real-world domains of learned behavior policies. However, given reliable and representative models, constraining physics parameters could be a good choice.
@@ 3.2 Training Data
We generate training light curves as follows. We place transients randomly in the 3-D volume of the Universe observable by ZTF. ZTF is a public-private partnership survey of the Northern sky conducted in $g$, $r$, and $i$ filters using a wide field (47 deg2) imager on the Palomar 48-inch telescope (Masci et al., 2019; Dekany et al., 2020; Graham et al., 2019; Bellm et al., 2019). For Type Ia supernovae we use SALT2 (Guy et al., 2007) models to estimate photometry in ZTF filters. We apply Milky way extinction (Cardelli et al., 1989; Schlegel et al., 1998) given the supernova sky location, redshift, and K-corrections. For core-collapse supernovae, we use 2-D Gaussian Process fits to ZTF light curves of spectroscopically confirmed supernovae of types II, IIn, IIb, Ib, Ic, Ic-BL, and superluminous supernovae types I and II. While we apply redshift corrections (scaling and stretch), due to the empirical nature of the fit, we are unable to apply extinction corrections. We estimate K-correction by estimating the Gaussian Process fit in the wavelength dimension at the redshifted mean wavelength of the ZTF filter at the sampled redshift. For GRB afterglows, we use the semi-analytic models of Ryan et al. (2020). Finally, we use the models from Bulla (2019) to simulate KNe from BNS and NSBH mergers. For KNe we use the interpolation scheme in Pang et al. (2023). For these, once again, while redshift scaling is accounted for, we do not account for extinction or K-correction.
@@
Figure 1: Learning curve of our KN follow-up agent Pythia (blue). Our agent achieves >3.5×>3.5$\times$> 3.5 × higher score than a random agent during training. The score variance (shaded blue region) is computed using a moving average with a length of 50 episodes. The mean random score is 6/9. Human references are shown as orange shaded regions.
@@ 3.3 RL Algorithm
We learn the optimal behavior policy using SARSA. SARSA is a simple on-policy learning algorithm, i.e. samples from the policy are used to update $Q$ (see Algorithm 2). We parametrize $Q=x(s,a)^{T}w$, where $x(s,a)$ is the state-action feature representation and $w$ is a set of linear weights to be learned. We choose a TD(0) target, i.e. $r_{t}+\gamma\hat{Q}(s_{t+1},a_{t+1};\hat{w})$, where $\hat{Q}$ is the running estimate. The policy is $\epsilon$-greedy: with probability $\epsilon$ the agent chooses a random action and otherwise chooses $\max_{a}\hat{Q}(s_{t},a_{t};\hat{w})$, the action with the maximum $\hat{Q}$-value. If the learning rate $\alpha_{t}$ satisfies the Robbins-Munro sequence and $\epsilon\to 0$ as $i\to\infty$ the agent is Greedy in the Limit of Infinite Exploration (GLIE) and is a sufficient condition for convergence in SARSA [^3] Two conditions to be GLIE: (1) we visit every state-action pair infinite times and (2) behavior policy converges to greedy policy. We parameterize $\epsilon=1/k^{n}$, where $n$ is a hyperparameter that controls how fast the agent becomes greedy. This parameterization allows the agent to get in a lot of exploration early during training and gradually forces it become more greedy. To ensure sample efficiency and propagate TD updates faster, we reuse episodes $m$ times, where we set $m=5$. Our agent learns online (i.e. it collects new experiences as opposed to episodes collected beforehand) in a simulated environment [^4] It is possible to train the agent online with real GW trigger scenarios but RL algorithms are notoriously sample inefficient and can take a lot of exploration to arrive at the optimal $Q$/policy..
@@
Our choice of linear VFA was motivated by the fact that the triad of non-linear function approximators, off-policy learning (like Q-learning), and bootstrapping (as in Temporal-difference learning) can cause the $Q$ function to diverge (Tsitsiklis & Van Roy, 1997). Though recent works have made significant progress to address this issue (e.g. prioritized replay (Schaul et al., 2015), asynchronous methods (Mnih et al., 2016), dueling DQN targets (Wang et al., 2016)) we adopt linear VFA and SARSA for our toy implementation to assess the feasibility of RL approaches to solve the KN follow-up problem and disambiguate limitations from convergence issues. An issue may be that the choice of linear VFA as hypothesis function class may not be a sufficiently rich representation of the true $Q$ function. This does indeed appear to be the case as we discuss later.